OpenStudy (anonymous):

. Given the reaction, Fe2O3 + 3 CO arrow 2 Fe + 3 CO2, if 12.0 moles of CO2 are made, then how many moles of CO were used? (Points : 3) 12.0 moles 6.0 moles 4.0 moles 3.0 moles

OpenStudy (aaronq):

Use the stoichiometric coefficients to find moles produced. Set up a ratio using the species of interest, like so: e.g. for a general reaction: \(\color{red}{a}A + \color{blue}{b}B\) \(\rightleftharpoons\) \( \color{green}{c}C\) where upper case are the species (A,B,C), and lower case (a,b,c) are the coefficients , \(\dfrac{n_A}{\color{red}{a}}=\dfrac{n_B}{\color{blue}{b}}=\dfrac{n_C}{\color{green}{c}}\) From here you can isolate what you need. For example: if you have 2 moles of B, how many moles of C can you produce? solve algebraically: \(\dfrac{2}{\color{blue}{b}}=\dfrac{n_C}{\color{green}{c}}\rightarrow n_C=\dfrac{2*\color{green}{c}}{\color{blue}{b}}\)

OpenStudy (anonymous):

\[ \frac{ 2c }{ b }\]

OpenStudy (anonymous):

I think its 4 moles

OpenStudy (anonymous):

@aaronq

OpenStudy (aaronq):

you would set up a ratio like this: \(\dfrac{n_{CO_2}}{3}=\dfrac{n_{CO}}{3}\) notice that the coefficients are the same, plug in your values \(\dfrac{12~~moles}{3}=\dfrac{n_{CO}}{3}\) without doing anything you can see that \(n_{CO}=12~moles\) because the coefficients are the same.